OpenStudy (anonymous):

-x+y=1 x+y=0 who know how to solve this problem and graph it i need helpppppppppppppppppp

OpenStudy (agent0smith):

Use elimination. Add them together, vertically, like so -x+y=1 + x+y=0 --------- 0+2y=1

OpenStudy (anonymous):

do you know how to graph it

OpenStudy (loser66):

do you know ? @jinan

OpenStudy (anonymous):

know what Loser66

OpenStudy (loser66):

graph them. They are 2 equations, right? so, we have to say "them"

OpenStudy (anonymous):

i think i have to solve the problem that graph it

OpenStudy (anonymous):

i dont know how to graph it

OpenStudy (loser66):

neither I, @agent0smith help us, please!!!

OpenStudy (agent0smith):

Do you know how to graph a line like y = mx + b?

OpenStudy (anonymous):

no

OpenStudy (agent0smith):

Okay well you'll need to learn how to do that. Try this https://www.youtube.com/watch?v=rgvysb9emcQ#at=60

OpenStudy (anonymous):

is hard

OpenStudy (agent0smith):

I understand, but it'll be much easier to learn from a video than me trying to teach you by typing everything out.